OpenStudy (anonymous):

1/256=16^x-9 Can someone please show me how to solve this?

OpenStudy (ash2326):

Is this your question \[\large\frac{1}{256}=16^{x-9}\]

OpenStudy (ash2326):

??

OpenStudy (anonymous):

Yes it is.

OpenStudy (ash2326):

Ok tell me what's the power of 2 so that we get 256 \[2^y=256\] or find y?

OpenStudy (anonymous):

1/256 = 16^(x -9) 1/4^4 = 4^2(x - 9) 1/4^4 can be written as 4^-4 using the laws of indices => 4^-4 = 4^2(x - 9) Since the bases are the same, they cancel out leaving: -4 = 2(x - 9) -4 = 2x - 18 -4 + 18 = 2x 14 = 2x x = 7

OpenStudy (anonymous):

Thank you so much!